Suppose that you are 20 years old today, and you want to have $500,000 when you reach age of 30. how much would you have to invest today to reach your goal? Suppose the interest rate is equal to 5%.
Answer #1

Using formula:
Present value=Future Value/(1+rate)^t=500000/1.05^(50-20)=115688.72433

Using financial calculator:
N=50-20=30
PMT=0
I/Y=5%
FV=-500000
CPT PV=115688.72433

Using excel:
=PV(5%,50-20,0,-500000)
=115688.72433